7. It is generally recommended and normal for a designer to add a hunting tooth to evenly distribute the wear on meshing teeth like a 53 teeth mesh with 30 teeth pinion. Find the rotation of pinion to mesh same pair of teeth again.


1
Expert's answer
2022-04-08T09:43:03-0400

So, number of rotation of pinion = LCM(53;30)/30 = 53×30/30 = 53

LCM - least common multiple